# V2X-DGW: Domain Generalization for Multi-agent Perception under Adverse Weather Conditions

### v5

@@ Abstract
Current LiDAR-based Vehicle-to-Everything (V2X) multi-agent perception systems have shown the significant success on 3D object detection. While these models perform well in the trained clean weather, they struggle in unseen adverse weather conditions with the domain gap. In this paper, we propose a Domain Generalization based approach, named V2X-DGW, for LiDAR-based 3D object detection on multi-agent perception system under adverse weather conditions. Our research aims to not only maintain favorable multi-agent performance in the clean weather but also promote the performance in the unseen adverse weather conditions by learning only on the clean weather data. To realize the Domain Generalization, we first introduce the Adaptive Weather Augmentation (AWA) to mimic the unseen adverse weather conditions, and then propose two alignments for generalizable representation learning: Trust-region Weather-invariant Alignment (TWA) and Agent-aware Contrastive Alignment (ACA). To evaluate this research, we add Fog, Rain, Snow conditions on two publicized multi-agent datasets based on physics-based models, resulting in two new datasets: OPV2V-w and V2XSet-w. Extensive experiments demonstrate that our V2X-DGW achieved significant improvements in the unseen adverse weathers.
@@ I Introduction
The advent of multi-agent perception systems has marked a significant leap forward in surmounting the inherent limitations of single-agent perception from the challenges of perceiving range and occlusion [1, 2]. This paper uses the Connected and Automated Vehicles (CAV) based cooperative perception in autonomous driving as an example of the multi-agent perception systems. Leveraging the Vehicle-to-Everything (V2X) communication for the exchange of LiDAR based point cloud data, these multi-agent systems have demonstrated a substantial enhancement in 3D object detection over the single-agent perception systems [3, 4, 5, 6]. However, the LiDAR-based V2X multi-agent perception systems under the adverse weather conditions (Fog, Rain, Snow) have not been studied and evaluated before.
@@
  • Accumulated Damages on Perception: Intuitively, the adverse weather conditions impair each agent’s perception, so the multi-agent perception system might have accumulated damages due to the agent-to-agent information sharing.
@@ II Related Work
Multi-agent Perception. Multi-agent perception systems transcend the constraints of single-vehicle systems by utilizing information via collaboration modules, enhancing effectiveness and results of multi-agent perception tasks [12, 4, 3, 8]. Generally, three methods are mainly used for combining observations from multiple vehicles: merging raw data [13], integrating features during processing [5, 1, 3, 4, 6], and output fusion [14, 15]. Attfuse [4] utilizes self-attention models to enhance the sharing features for V2V perception system. SCOPE [16] presents a learning-based framework that targets multi-agent challenges, prioritizing the temporal context of the ego agent. V2X-ViT [8] introduces a unified Vision Transformer architecture to integrate features from multi-agent perception systems. While these methods demonstrate remarkable performance in multi-agent perception, they have primarily been evaluated under clear weather conditions, neglecting the impact of adverse weather.
Domain Generalization.  Domain generalization (DG) aims at achieving the excellent performance in unseen domains by exclusively learning from source domains. Current DG methods have been extensively explored on object detection [17, 18], semantic segmentation [19, 20, 21, 22] and reinforcement learning [23, 24] tasks in 2D and 3D computer vision tasks. As for strategies of DG employing multiple sources, which encompasses techniques such as the the disentanglement of domain-specific and domain-invariant features [25, 26, 27], alignment of source domain distributions [28, 29, 30], and domain augmentation [31]. These approaches have been instrumental in navigating the complexities associated with learning from diverse source domains. Some recent studies [32, 22, 21] have analyzed the sparsity of point clouds in addressing DG issues. However, the extension of DG to the LiDAR-based 3D object detection on multi-agent perception systems in adverse weather conditions remains relatively untapped. This paper aims to design a new DG method to promote the LiDAR-based 3D object detection on multi-agent perception systems in adverse weather conditions.
@@ III-B Overview of Architecture
Figure 3: Illustration of Adaptive Weather Augmentation.
@@ IV-B Quantitative Evaluation
Figure 6: Sensitivity to agent numbers. Other methods’ decreasing detection performance with more agent numbers on OPV2V-w indicates the accumulated damages under adverse weather, while our method is more robust.
Sensitivity to Agent Numbers. The Fig. 6 illustrates the impact of cooperative perception performance with varying numbers of agents under rain and snow weather. Intuitively, the cooperative perception performance should be improved with an increasing number of nearby agents [4]. However, most comparison methods sometimes have decreasing cooperative detection performance with increasing agent numbers under adverse weather scenarios, which is because of their sensitivity to the accumulated damages from more agents in adverse weather. Differently, our V2X-DGW could overcome the challenge of accumulated damages.

### v6

@@ Abstract
Current LiDAR-based Vehicle-to-Everything (V2X) multi-agent perception systems have shown the significant success on 3D object detection. While these models perform well in the trained clean weather, they struggle in unseen adverse weather conditions with the domain gap. In this paper, we propose a Domain Generalization based approach, named V2X-DGW, for LiDAR-based 3D object detection on multi-agent perception system under adverse weather conditions. Our research aims to not only maintain favorable multi-agent performance in the clean weather but also promote the performance in the unseen adverse weather conditions by learning only on the clean weather data. To realize the Domain Generalization, we first introduce the Adaptive Weather Augmentation (AWA) to mimic the unseen adverse weather conditions, and then propose two alignments for generalizable representation learning: Trust-region Weather-invariant Alignment (TWA) and Agent-aware Contrastive Alignment (ACA). To evaluate this research, we add Fog, Rain, Snow conditions on two publicized multi-agent datasets based on physics-based models, resulting in two new datasets: OPV2V-w and V2XSet-w. Extensive experiments demonstrate that our V2X-DGW achieved significant improvements in the unseen adverse weathers. The code is available at https://github.com/Baolu1998/V2X-DGW.
@@ I Introduction
The advent of multi-agent perception systems has marked a significant leap forward in surmounting the inherent limitations of single-agent perception from the challenges of perceiving range and occlusion [1, 2]. This paper uses the Connected and Automated Vehicles (CAV) based cooperative perception in autonomous driving as an example of the multi-agent perception systems. Leveraging the Vehicle-to-Everything (V2X) communication for the exchange of LiDAR based point cloud data, these multi-agent systems have demonstrated a substantial enhancement in 3D object detection over the single-agent perception systems [3, 4, 5, 6, 7, 8, 9, 10]. However, the LiDAR-based V2X multi-agent perception systems under the adverse weather conditions (Fog, Rain, Snow) [11] have not been studied and evaluated before.
@@
  • Accumulated Damages on Perception: Intuitively, the adverse weather conditions impair each agent’s perception, so the multi-agent perception system might have accumulated damages [13] due to the agent-to-agent information sharing.
@@ II Related Work
Multi-agent Perception. Multi-agent perception systems transcend the constraints of single-vehicle systems by utilizing information via collaboration modules, enhancing effectiveness and results of multi-agent perception tasks [18, 4, 3, 14, 19]. Generally, three methods are mainly used for combining observations from multiple vehicles: merging raw data [20], integrating features during processing [5, 1, 3, 4, 6], and output fusion [21, 22]. Attfuse [4] utilizes self-attention models to enhance the sharing features for V2V perception system. SCOPE [23] presents a learning-based framework that targets multi-agent challenges, prioritizing the temporal context of the ego agent. V2X-ViT [14, 7] introduces a unified Vision Transformer architecture to integrate features from multi-agent perception systems. While these methods demonstrate remarkable performance in multi-agent perception, they have primarily been evaluated under clear weather conditions, neglecting the impact of adverse weather.
Domain Generalization.  Domain generalization (DG) aims at achieving the excellent performance in unseen domains by exclusively learning from source domains. Current DG methods have been extensively explored on object detection [24, 25], semantic segmentation [26, 27, 28, 29, 30] and reinforcement learning [31, 32] tasks in 2D and 3D computer vision tasks. As for strategies of DG employing multiple sources, which encompasses techniques such as the the disentanglement of domain-specific and domain-invariant features [33, 34, 35], alignment of source domain distributions [36, 37, 38], and domain augmentation [39]. These approaches have been instrumental in navigating the complexities associated with learning from diverse source domains. Some recent studies [40, 29, 28] have analyzed the sparsity of point clouds in addressing DG issues. However, the extension of DG to the LiDAR-based 3D object detection on multi-agent perception systems in adverse weather conditions remains relatively untapped. This paper aims to design a new DG method to promote the LiDAR-based 3D object detection on multi-agent perception systems in adverse weather conditions.
@@ III-B Overview of Architecture
Figure 3: Illustration of simulation before and after Adaptive Weather Augmentation.
@@ IV-B Quantitative Evaluation
Figure 6: Sensitivity to agent numbers. Other methods’ decreasing detection performance with more agent numbers on OPV2V-w (contrary to intuition [4]) indicates the accumulated damages under adverse weather, while our method is more robust.
Sensitivity to Agent Numbers. The Fig. 6 illustrates the impact of cooperative perception performance with varying numbers of agents under rain and snow weather. As shown in the intuition [4], the cooperative perception performance should be improved with an increasing number of nearby agents. However, most comparison methods sometimes have decreasing cooperative detection performance with increasing agent numbers under adverse weather scenarios, which is because of their sensitivity to the accumulated damages from more agents in adverse weather. Differently, our V2X-DGW could overcome the challenge of accumulated damages.